OpenStudy (anonymous):

find the quotient. -300x^5y^4 % xy^2 % means the divide sign:P

OpenStudy (anonymous):

\[\frac{ x^a }{ x^b } = x^{a-b}\] \[-\frac{ 300x^5y^4 }{ xy^2 } = -300 \times \frac{ x^5 }{ x } \times \frac{ y^4 }{ y^2 } = -300x^{5-1}y^{4-2} = -300x^4y^2\] also, you'll want to know that:\[x^a \times x^b = x^{a+b}\]and\[(x^a)^b = x^{ab} = (x^b)^a\]

OpenStudy (anonymous):

thank you so much